OpenStudy (anonymous):

for f(x)=3x+1 and g(x)=x^2-6 find (f+g)(x)

OpenStudy (trojanpoem):

All you have to do , is to apply the operation it asks for : (f+g)(x) = f(x) + g(x) = [3x +1 ] + [x^2 -6] = x^2 + 3x - 5

OpenStudy (anonymous):

thx